if a tank holds 100Gallons with a mix of 60% water and 40%oil , how many gallons of oil does tank contain?
Answer by Alan3354(69443)

if a tank holds 100Gallons with a mix of 60% water and 40%oil , how many gallons of oil does tank contain?
------
40% of 100 gal = (40/100)*100 = 0.4*100 = 40 gallons